OpenStudy (anonymous):

Of the 18 coins in my pocket, I had a some pennies, twice as many twenty pence pieces and the rest fifty-pence pieces. If I had 4.64 in total, how many of each type of coin did I have.

OpenStudy (chris):

There are there unknown variables here, so you're going to need three equations. Do you know any of the equations that the word problem describes?

OpenStudy (chris):

Let's call the three unknowns "P" for penny, "T" for twenty, and "F" for fifty.

OpenStudy (anonymous):

What is the relation between Penny, Twenty and fifty?

OpenStudy (chris):

So one equation is that there are 18 coins: 1. P + T + F = 18

OpenStudy (anonymous):

\( T=2P \) and \( F = 18-3P \)

OpenStudy (anonymous):

using Chris notation.

OpenStudy (chris):

=) So we need just one more.

OpenStudy (anonymous):

\( Px+Ty+Fz = 4.64 \)

OpenStudy (anonymous):

I don't know your system, so i can't assign values to the x,y and z :(

OpenStudy (chris):

ah, I was thinking P + 20T + 50F = 4.64 (should have lowercased my variables)

OpenStudy (chris):

heh my bad, by P/T/F I meant the variable names, so p/t/f

OpenStudy (anonymous):

4.64 is \( 4.64\$ \) ?

OpenStudy (anonymous):

I got it!

OpenStudy (chris):

ah good question though

OpenStudy (anonymous):

@Chris: If penny is P then twenty pence is 20 P ?

OpenStudy (chris):

I'm assuming that it's actually .01p + .20t + .50f = 4.64

OpenStudy (anonymous):

and 1 penny is 1/100th of a dollar?

OpenStudy (chris):

correct

OpenStudy (chris):

well, Australian dollars

OpenStudy (chris):

1 pence = 1 dollar/100

OpenStudy (anonymous):

lol, I think they should give this information with the question itself.

OpenStudy (anonymous):

Its a british textbook, so I guess we are dealing with £££

OpenStudy (anonymous):

this question is easy if you know this information, and if you don't it's impossible to solve with certainty.

OpenStudy (chris):

haha, so pence = 1 pound/100 right?

OpenStudy (chris):

I think we have it right ffm

OpenStudy (anonymous):

Yes I think that too chris.

OpenStudy (chris):

three equations: 1. p + t + f = 18 2. p + .2t + .5f = 4.64 3. t = 2p Where p = number of pennies, t = number of twenty pence, and f = number of fifty pence

OpenStudy (anonymous):

Now with this three equation, you could just use substitution and get your desired result

OpenStudy (chris):

You know how to do that chrissy?

OpenStudy (anonymous):

i remember, finding unknown variables using 3 equations

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

You guys have me very helpful

OpenStudy (anonymous):

*been

OpenStudy (anonymous):

since we alreeady have the value of T=2p, substitute to the 2 equation then 2 equations will remain, you can already solve system of 2 equation using elimination and substitution, choose. .

OpenStudy (chris):

No problem! I think the trickiest part of this problem was that line - "twice as many twenty pence", if you skip over that then you can never get the third equation. t = 2p.

OpenStudy (anonymous):

@jerwyn gayo: There are different ways that we could use to solve a system of three equations with three unknowns for example you could use Matrix algebra, but for this purpose the best way is to use substitution.

OpenStudy (anonymous):

@ffm. yeah,.